OpenStudy (mellyg):

Can somebody help me on how to find the maximum and minimums of quadratic functions? Thanks so much.

OpenStudy (anonymous):

are you in or have you taken calculus yet?

OpenStudy (mellyg):

no. im in algebra 1

OpenStudy (anonymous):

ok, I forgot how to do it by algebra, sorry I can't help

OpenStudy (mellyg):

its fine, thanks so much

OpenStudy (arthur326):

The standard method is completing the square. For example, suppose you have the quadratic \[x^2-2x+4.\] You can rewrite this as \[(x^2-2x + 1) + 3 = (x-1)^2 +3.\] Now we see that \[(x-1)^2\] is always non-negative, so this particular quadratic has a minimum at \[x=1,\] and it has a value of 3 there. If you do this process in general, you will find that the quadratic \[ax^2+bx+c\] has a minimum/maximum at \[x=-\frac{b}{2a}.\] (This is where the vertex of the parabola is if you graph the equation.)

OpenStudy (mellyg):

how do you complete the square? Im confused on that part

OpenStudy (arthur326):

First suppose you're given a quadratic in the form \[x^2+ax+b.\] We want to rewrite this as \[(x+r)^2\] for some r. Now we know that \[(x+r)^2 = x^2+2rx+r^2.\] So we set \[2r=a \quad \Rightarrow \quad r=\frac{a}{2}.\] But \[(x+a/2)^2 = x^2 + ax + \frac{a^2}{4},\] so we need to add \[b^2-\frac{a^2}{4}\] to both sides to get the original quadratic. So after completing the square our quadratic becomes \[(x+a/2)^2 + b^2-\frac{a^2}{4}.\] Now if you have a quadratic of the form \[ax^2+bx+c\] (where the coefficient of the square term doesn't have to be 1), just factor a from the expression and proceed in the same way for the inside quadratic.

OpenStudy (mellyg):

what does the r represent?

OpenStudy (arthur326):

It's just a temporary variable I made to represent a constant that is added to x before we square the expression. That was just the form we wanted to rewrite the quadratic in order to complete the square. The first quadratic reaches its minimum when \[x=-r\] because that makes the squared term 0. Now I should have added that we were trying to rewrite the quadratic \[x^2+ax+b\] in the form \[(x+r)^2 + k\] for some constant k. In the end that's what we did.

OpenStudy (mellyg):

the function that i am trying to do this on is f(x)= 3x^2+6x+7. Could you walk me through the problem using these formulas? Im still kind of confused on what the variables represent in the equation

OpenStudy (arthur326):

Okay. Let's factor out 3 from the first two terms: \[3(x^2+2x)+7.\] Now I want to make \[(x^2+2x)\] a perfect square by adding a constant to it: \[3(x^2+2x+t)+7.\] However, that just increased the function by 3t (note that there's a 3 outside.) So in order for the expression to remain the same, I have to subtract 3t at the end: \[3(x^2+2x+t)+7 -3t.\] Now we have to choose a t such that \[(x^2+2x+t)\] is a perfect square. What t should we use?

OpenStudy (arthur326):

Hint: What is \[(x+1)^2?\]

OpenStudy (mellyg):

im not really sure what that means, but (x+1)^2 can be written as (x+1) (x+1), but if t is a perfect square (it has to be to be in a perfect square equation) than it could be something's square root? Im not sure but thats all i got lol

OpenStudy (arthur326):

Well t does not always have to be a perfect square, but you're right that it is in this case. Now you said that \[(x+1)^2=(x+1)(x+1),\] and that's right. But can you expand the answer further?

OpenStudy (mellyg):

not that i have learned, but whether you can or not is open for discussion

OpenStudy (johnweldon1993):

Woh, this whole process wasn't really needed! It's good to know but... To find the Maximum or Minimum point of a quadratic *also called the vertex*..use the equation that was posted in the very beginning Your function looks like \(\large y = ax^2 + bx + c\) The x-coordinate of the vertex can be found by using the formula \[\large x = -\frac{b}{2a}\] To then find the y-coordinate, simply plug in this 'x' you just found into the original function

OpenStudy (mellyg):

@johnweldon1993 , i got (-1,5), is this correct. If so, is it the minimum?

OpenStudy (mellyg):

lol wait i think its (-1,4)

OpenStudy (johnweldon1993):

\[\large y= 3x^2 + 6x+ 7\] So a = 3 b = 6 c = 7 \[\large x = -\frac{b}{2a} = -\frac{6}{2(3)} = -\frac{6}{6} = -1\] Good...now plugging that back in you get \[\large y(-1) = 3(-1)^2 + 6(-1) + 7\] \[\large y(-1) = 3(1) - 6 + 7\] \[\large y(-1) = 3 + 1\] \[\large y(-1) = 4\] So I have the vertex at (-1,4) indeed And that is a minimum you say? How do you know? :)

OpenStudy (mellyg):

because when you graph it, the lowest point of the parabola is the vertex (-1,4)?

OpenStudy (mellyg):

^ @johnweldon1993

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (mellyg):

thank you so much for your help! :) @johnweldon1993

OpenStudy (johnweldon1993):

Of course :)